### v1

# Skagit Fisheries Enhancement Group   UMass Amherst

Abstract
Object detectors often perform poorly on data that differs from their training set. Domain adaptive object detection (DAOD) methods have recently demonstrated strong results on addressing this challenge. Unfortunately, we identify systemic benchmarking pitfalls that call past results into question and hamper further progress: (a) Overestimation of performance due to underpowered baselines, (b) Inconsistent implementation practices preventing transparent comparisons of methods, and (c) Lack of generality due to outdated backbones and lack of diversity in benchmarks. We address these problems by introducing: (1) A unified benchmarking and implementation framework, Align and Distill (ALDI), enabling comparison of DAOD methods and supporting future development, (2) A fair and modern training and evaluation protocol for DAOD that addresses benchmarking pitfalls, (3) A new DAOD benchmark dataset, CFC-DAOD, enabling evaluation on diverse real-world data, and (4) A new method, ALDI++, that achieves state-of-the-art results by a large margin. ALDI++ outperforms the previous state-of-the-art by +3.5 AP50 on Cityscapes $\rightarrow$ Foggy Cityscapes, +5.7 AP50 on Sim10k $\rightarrow$ Cityscapes (where ours is the only method to outperform a fair baseline), and +2.0 AP50 on CFC Kenai $\rightarrow$ Channel. Our framework777github.com/justinkay/aldi, dataset888github.com/visipedia/caltech-fish-counting, and state-of-the-art method offer a critical reset for DAOD and provide a strong foundation for future research.
Keywords:
Domain adaptation Object detection
1 Introduction
Figure 1: ALDI achieves state-of-the-art performance in domain adaptive object detection (DAOD) and provides a unified framework for fair comparison. We show: (1) Inconsistent implementation practices give the appearance of steady progress in DAOD (left bars [12, 10, 8, 26, 34]); reimplementation and fair comparison with ALDI shows less difference between methods than previously reported (middle bars); (2) A fairly constructed source-only model (blue line) outperforms many existing DAOD methods, indicating less progress has been made than previously reported; and a proper oracle (orange line) outperforms all existing methods, in contrast to previously-published results; and (3) Our proposed method ALDI++ (green bars) achieves state-of-the-art performance on DAOD benchmarks such as Cityscapes normal-$\rightarrow$ Foggy Cityscapes and is complementary to ongoing advances in object detection like VitDet [33].
The challenge of DAOD. Modern object detector performance, though excellent across many benchmarks [36, 53, 52, 3, 49, 46], often severely degrades when test data exhibits a distribution shift with respect to training data [41]. For instance, detectors do not generalize well when deployed in new environments in environmental monitoring applications [30, 53]. Similarly, models in medical applications perform poorly when deployed in different hospitals or on different hardware than they were trained [55, 19]. Unfortunately, in real-world applications it is often difficult, expensive, or time-consuming to collect the additional annotations needed to address such distribution shifts in a supervised manner. An appealing option in these scenarios is unsupervised domain adaptive object detection (DAOD), which attempts to improve detection performance when moving from a “source” domain (used for training) to a “target” domain (used for testing) [32, 29] without the use of target-domain supervision.
The current paradigm. The research community has established a set of standard benchmark datasets and methodologies that capture the deployment challenges motivating DAOD. Benchmarks consist of labeled data that is divided into two sets: a source and a target, each originating from different domains. DAOD methods are trained with source-domain images and labels, as in traditional supervised learning, and have access to unlabeled target domain images. The target-domain labels are not available for training.
To measure DAOD methods’ performance, researchers use source-only models and oracle models as points of reference. Source-only models—sometimes also referred to as baselines—are trained with source-domain data only, representing a lower bound for performance without domain adaptation. Oracle models are trained with supervised target-domain data, representing a fully-supervised upper bound. The goal in DAOD is to close the gap between source-only and oracle performance without target-domain supervision.
Impediments to progress. Recently-published results indicate DAOD is exceptionally effective, doubling the performance of source-only models and even outperforming fully-supervised oracles [34, 8, 5]. However, upon close examination we discover problems with current benchmarking practices that call these results into question:
P1: Improperly constructed source-only and oracle models, leading to overestimation of performance gains. We find that source-only and oracle models are consistently constructed in a way that does not properly isolate domain-adaptation-specific components, leading to misattribution of performance improvements. We show that when source-only and oracle models are fairly constructed—i.e. use the same architecture and training settings as DAOD methods—no existing methods outperform oracles and many methods do not even outperform source-only models (Fig. 1), in stark contrast to claims made by recent work. These results mean we do not have an accurate measure of the efficacy of DAOD.
P2: Inconsistent implementation practices preventing transparent comparisons of methods. We find existing DAOD methods are built using a variety of different object detection libraries with inconsistent training settings, making it difficult to determine whether performance improvements come from new DAOD methods or simply improved hyperparameters. We find that tweaking these hyperparameters—whose values often differ between methods yet are not reported in papers—can lead to a larger change in performance than the proposed methods themselves (see Sec. 6.3), thus we cannot take reported advancements at face value. Without the ability to make fair comparisons we cannot transparently evaluate contributions nor make principled methodological progress.
P3: (a) Lack of diverse benchmarks and (b) outdated model architectures, leading to overestimation of methods’ generality. DAOD benchmarks have focused largely on urban driving scenarios with synthetic distribution shifts [48, 28], and methods continue to use outdated detector architectures for comparison with prior work [9]. The underlying assumption is that methods will perform equivalently across application domains and backbone architectures. We show that in fact the ranking of methods changes across benchmarks and architectures, revealing that published results may be uninformative for practitioners using modern architectures and real-world data.
A critical reset for DAOD research. DAOD has the potential for impact in a range of real-world applications, but these systemic benchmarking pitfalls impede progress. We aim to address these problems and lay a solid foundation for future progress in DAOD with the following contributions:
1. Align and Distill (ALDI), a unified benchmarking and implementation framework for DAOD. In order to enable fair comparisons, we first identify key themes in prior work (Sec. 2) and unify common components into a single state-of-the-art framework, ALDI (Sec. 3). ALDI facilitates detailed study of prior art and streamlined implementation of new methods, supporting future research.
2. A fair and modern training protocol for DAOD methods, enabled by ALDI. We provide quantitative evidence of the benchmarking pitfalls we identify and propose an updated training and evaluation protocol to address them (Sec. 6.1). This enables us to set more realistic and challenging targets for the DAOD community and perform the first fair comparison of prior work in DAOD (Sec. 6.2).
3. A new benchmark dataset, CFC-DAOD, sourced from a real-world adaptation challenge in environmental monitoring (Sec. 5). CFC-DAOD increases the diversity of DAOD benchmarks and is notably larger than existing options. We show that the ranking of methods changes across different benchmarks (Sec. 6.2), thus the community will benefit from an additional point of comparison.
4. A new method, ALDI++, that achieves state-of-the-art results by a large margin. Using the same model settings across all benchmarks, ALDI++ outperforms the previous state-of-the-art by +3.5 AP50 on Cityscapes $\rightarrow$ Foggy Cityscapes, +5.7 AP50 on Sim10k $\rightarrow$ Cityscapes (where ours is the only method to outperform a fair source-only model), and +2.0 AP50 on CFC Kenai $\rightarrow$ Channel.
2 Related Work
Two methodological themes have dominated recent DAOD research: feature alignment and self-training/self-distillation. We first give an overview of these themes and previous efforts to combine them, and then use commonalities to motivate our unified framework, Align and Distill, in Sec. 3.
Feature alignment in DAOD. Feature alignment methods aim to make target-domain data “look like” source-domain data, reducing the magnitude of the distribution shift. The most common approach utilizes an adversarial learning objective to align the feature spaces of source and target data [16, 10, 9, 57]. Faster R-CNN in the Wild [9] utilizes adversarial networks at the image and instance level. SADA [10] extends this to multiple adversarial networks at different feature levels. Other approaches propose mining for discriminative regions [57], weighting local and global features differently [47], incorporating uncertainty [40], and using attention networks [51]. Alignment at the pixel level has also been proposed using image-to-image translation techniques to modify input images directly [12].
Self-training/self-distillation in DAOD. Self-training methods use a “teacher” model to predict pseudo-labels on target-domain data that are then used as training targets for a “student” model. Self-training can be seen as a type of self-distillation [43, 6], which is a special case of knowledge distillation [25, 7] where the teacher and student models share the same architecture. Most recent self-training approaches in DAOD are based on the Mean Teacher [50] framework, in which the teacher model is updated as an exponential moving average (EMA) of the student model’s parameters. Extensions to Mean Teacher for DAOD include: MTOR, which utilizes graph structure to enforce student-teacher feature consistency [4], Probabilistic Teacher (PT), which uses probabilistic localization prediction and soft distillation losses [8], and Contrastive Mean Teacher (CMT), which uses MoCo [21] to enforce student-teacher feature consistency [5].
Combining feature alignment and self-training. Several approaches utilize both feature alignment and self-training/self-distillation, motivating our unified framework. Unbiased Mean Teacher (UMT) [12] uses mean teacher in combination with image-to-image translation to align source and target data at the pixel level. Adaptive Teacher (AT) [55] uses mean teacher with an image-level discriminator network. Masked Image Consistency (MIC) [26] uses mean teacher, SADA, and a masking augmentation to enforce teacher-student consistency. Because these methods were implemented in different codebases using different training recipes and hyparameter settings, it is unclear which contributions are most effective and to what extent feature alignment and self-training are complementary. We address these issues by reimplementing these approaches in the ALDI framework and perform fair comparisons and ablation studies in Sec. 6.
DAOD datasets. Cityscapes (CS) $\rightarrow$ Foggy Cityscapes (FCS) [11, 48] is a popular DAOD benchmark that emulates domain shift caused by changes in weather in urban driving scenarios. The dataset contains eight vehicle and person classes. Sim10k $\rightarrow$ CS [28] poses a Sim2Real challenge, adapting from video game imagery to real-world imagery. The benchmark focuses on a single class, “car”. Other common tasks include adapting from real imagery in PascalVOC [15] to clip art and watercolor imagery [27]. We report results on CS $\rightarrow$ FCS and Sim10k $\rightarrow$ CS due to their widespread popularity in the DAOD literature and focus on real applications. We note that existing benchmarks reflect a relatively narrow set of potential DAOD applications. To study whether methods generalize outside of urban driving scenarios, in Sec. 5 we introduce a novel dataset sourced from a real-world adaptation challenge in environmental monitoring, where imagery is much different from existing benchmarks.
3 Align and Distill (ALDI): Unifying DAOD
We first introduce Align and Distill (ALDI), a new benchmarking and implementation framework for DAOD. ALDI unifies existing approaches in a common framework, enabling fair comparisons and addressing P2. Inconsistent implementation practices, while also providing the foundation for development of a new method ALDI++ that achieves state-the-art performance (Sec. 4, Sec. 6.2). The framework is visualized in Fig. 2. All components are ablated in Sec. 6.3.
Data. DAOD involves two datasets: a labeled source dataset $X_{src}$ and an unlabeled target dataset $X_{tgt}$. Each training step, a minibatch of size $B$ is constructed containing both $B_{src}$ source images and $B_{tgt}$ target images.
Models. A student model $\theta_{stu}$ and a teacher model $\theta_{tch}$ are initialized with the same weights, typically obtained through supervised pretraining on ImageNet, COCO, or $X_{src}$. Pretraining on $X_{src}$ is often referred to as “burn-in.” The student is trained through backpropagation. The teacher’s weights are not updated by backpropagation, rather they are updated to be the EMA of student weights [50].
Training involves three objectives:
1. Supervised training with source data. Each labeled source sample $x_{src,i}$ is transformed by some $t\sim T_{src}$, the set of possible source-domain transformations, then passed through the student model to obtain a supervised loss $L_{sup}$ given ground truth targets $y_{src,i}$. In our case $L_{sup}$ are Faster R-CNN losses [44].
2. Self-distillation with target data. Each unlabeled target sample $x_{tgt,i}$ passes through both the teacher and student models. The teacher’s predictions $\hat{p}_{tgt,i}$ act as distillation targets for the student’s predictions $p_{tgt,i}$, resulting in distillation losses $L_{distill}$ that are backpropagated through the student. Before computing $L_{distill}$ the teacher’s outputs are postprocessed to be either soft (e.g. logits or softmax outputs) or hard (e.g. thresholded pseudo-label) targets.
Before passing through the teacher, $x_{tgt,i}$ is transformed by some $\hat{t}\sim\hat{T}_{weak}$, a set of “weak” transformations—e.g. random horizontal flipping—that allow the teacher to provide high quality predictions $\hat{p}_{tgt}$. The same image is also passed to the student, this time transformed by $t\sim T_{tgt}$ which typically contains “stronger” augmentations such as color jitter or random erasing.
Table 1: ALDI unifies and extends existing work. We show settings to reproduce five prior works and our method ALDI++. See Fig. 2 and Sec. 3 for more details. Burn-in: fixed duration (Fixed), our approach (Ours, Sec. 4). Augs. $T_{src},T_{tgt}$: Random flip (F), multi-scale (M), crop & pad (CP), color jitter (J), cutout [13] (C), MIC [26]. ${\frac{1}{2}}$: augs used on half the images in the batch. $\frac{B_{tgt}}{B}$: Target-domain portion of minibatch of size $B$. Postprocess: Processing of teacher preds before distillation: sigmoid/softmax (Sharpen), sum class preds for pseudo-objectness (Sum), conf. thresholding (Thresh), NMS. $L_{distill}$: Distillation losses: hard pseudo-labels (Hard), continuous targets (Soft). $L_{align}$: Feature alignment losses: image-level adversarial (Img), instance-level adversarial (Inst), image-to-image translation (Img2Img). ${}^{\dagger}$: settings used in ALDI implementation (last column) but not in the original implementation (second-to-last column). ${}^{at}$: source-only and oracle results sourced from [34].
[…]
Burn-in $T_{src}$ $T_{tgt}$ $\frac{B_{tgt}}{B}$ process $L_{distill}$ $L_{align}$ (Reported) (w/ ALDI)
Source-only F, ${}^{\dagger}$, ${}^{\dagger}$, ${}^{\dagger}$ 0.0 ${}^{at}$23.5 59.1
SADA [10] F F 0.5 Img, Inst 44.0 54.2
PT [8] Fixed F F, J, C 0.3 Sharpen, Sum Soft 47.1 59.2
UMT [12] CP, J 0.5 Thresh, NMS Hard Img2Img 41.7 61.4
MIC [26] F F, J, MIC 0.5 Thresh, NMS Hard Img, Inst 47.6 61.7
[…]
ALDI++ Ours F, M, J, C F, M, J, MIC 0.5 Sharpen Soft 66.8
Oracle F, ${}^{\dagger}$, ${}^{\dagger}$, ${}^{\dagger}$ 1.0 ${}^{at}$42.7 67.2
Figure 2: (Left) The ALDI student-teacher framework for DAOD. Each training step (moving left to right and bottom to top): (1) Sample $B_{src}$ labeled source images $x_{src}$; transform by $t\sim T_{src}$; pass to student; compute supervised loss $L_{sup}$ using ground-truth labels $y_{src}$. (2) Sample $B_{tgt}$ unlabeled target images $x_{tgt}$; transform by $t\sim T_{tgt}$; pass to student to get preds $p_{tgt}$. Compute alignment objectives $L_{align}$ using $x_{src}$ and $x_{tgt}$. (3) Pass same unlabeled target data $x_{tgt}$, weakly transformed, to teacher; postprocess to obtain teacher preds $\hat{p}_{tgt}$. Compute distillation loss $L_{distill}$ between teacher and student preds. Use stop gradient (SG) on teacher model; update teacher to the EMA of student’s weights. (Middle, Right) ALDI++ (Sec. 4) introduces two new methods to achieve state-of-the-art performance: (Middle) A robust burn-in strategy utilizing strong augmentations and EMA, and (Right) Multi-task soft distillation losses to train the student using teacher outputs at all detector stages. $\sigma$: sigmoid or softmax for binary cross-entropy and cross entropy losses, respectively.
3. Feature alignment. $x_{src,:}$ and $x_{tgt,:}$ are “aligned” via an alignment objective $L_{align}$ that enforces invariance across domains either at the image or feature level.
Unification of prior work. We demonstrate the generality of our framework by reimplementing five recently-proposed methods on top of ALDI for fair comparison: UMT [12], SADA [10], PT [8], MIC [26], and AT [34]. In Tab. 1 we enumerate the settings required to reproduce each method.
4 ALDI++: Improving DAOD
We next propose two novel enhancements to the Align and Distill approach, resulting in a new method ALDI++. We show in Sec. 6.2 that these enhancements lead to state-of-the-art results, and ablate each component in Sec. 6.3.
1. Robust burn-in. First we propose a new “burn-in” strategy for pretraining a teacher model $\theta_{tch}$ on source-only data $X_{src}$. A key challenge in student-teacher methods is improving target-domain pseudo-label quality. We point out that psuedo-label quality in the early stages of self-training is largely determined by the out-of-distribution (OOD) generalization capabilities of the initial teacher model $\theta^{init}_{tch}$, and thus propose a training strategy aimed at improving OOD generalization during burn-in. We add strong data augmentations including random resizing, color jitter, and random erasing, and keep an EMA copy of the model during burn-in, two strategies that have previously been shown to improve OOD generalization and robustness [1, 2, 18]. We are the first to utilize these strategies for DAOD burn-in.
2. Multi-task soft distillation. Most prior work utilizes confidence thresholding and non-maximum suppression to generate “hard” pseudo-labels from teacher predictions $\hat{p}_{tgt}$ (see Tab. 1). However in object detection this strategy is sensitive to the confidence threshold chosen, leading to both false positive and false negative errors that harm self-training [31]. We take inspiration from the knowledge distillation literature and propose instead using “soft” distillation losses—i.e. using teacher prediction scores as targets without thresholding—allowing us to eliminate the confidence threshold hyperparameter.
We distill each task of Faster R-CNN—Region Proposal Network localization ($rpn$) and objectness ($obj$), and Region-of-Interest Heads localization ($roih$) and classification ($cls$)—independently. At each stage, the teacher provides distillation targets for the same set of input proposals used by the student—i.e. anchors $A$ in the first stage, and student region proposals $p^{rpn}_{tgt}$ in the second stage: $p^{rpn,obj}_{tgt}=\theta^{rpn,obj}_{stu}(A,x^{t}_{tgt})$ (1) $\hat{p}^{rpn,obj}_{tgt}=\theta^{rpn,obj}_{tch}(A,x^{\hat{t}}_{tgt})$ (2) $p^{roih,cls}_{tgt}=\theta^{roih,cls}_{stu}(p^{rpn}_{tgt},x^{t}_{tgt})$ (3) $\hat{p}^{roih,cls}_{tgt}=\theta^{roih,cls}_{tch}(p^{rpn}_{tgt},x^{\hat{t}}_{tgt})$ (4)
At each iteration, student distillation losses $L_{distill}$ are computed as:
$L^{rpn}_{distill}=\lambda_{0}L_{rpn}(p^{rpn}_{tgt},\hat{p}^{rpn}_{tgt})+\lambda_{1}L_{obj}(p^{obj}_{tgt},\hat{p}^{obj}_{tgt})$ (5)
$L^{roih}_{distill}=\lambda_{2}L_{roih}(p^{roih}_{tgt},\hat{p}^{roih}_{tgt})+\lambda_{3}L_{cls}(p^{cls}_{tgt},\hat{p}^{cls}_{tgt})$ (6)
$L_{distill}=L^{rpn}_{distill}+L^{roih}_{distill}$ (7)
Where $L_{rpn}$ and $L_{roih}$ are the smooth L1 loss and $L_{obj}$ and $L_{cls}$ are the cross-entropy loss, and $\lambda_{0\dots 3}=1$ by default. See Fig. 2 for a visual depiction. We include more implementation details in the supplemental material.
One prior DAOD work, PT [8], has also used soft distillation losses, however we note two shortcomings our method addresses: (1) PT requires a custom “Probabilistic R-CNN” architecture for distillation, while our approach is general and can work with any two-stage detector, and (2) PT uses $\hat{p}^{cls}$ as an indirect proxy for distilling $p^{obj}$, while our approach is able to distill each task directly.
5 The CFC-DAOD Dataset
Next we introduce our dataset contribution, CFC-DAOD, addressing P3: (a) Lack of diverse benchmarks leading to overestimation of methods’ generality.
Figure 3: The CFC-DAOD benchmark focuses on detecting fish (white bounding boxes) in sonar imagery under domain shift caused by environmental differences between the training location (Kenai) and testing location (Channel). Our dataset contains 168k bounding boxes in 29k frames sampled from 150 new videos captured over two days from 3 different sonar cameras on the Channel river, enabling DAOD experiments. Here we visualize the distribution of images and annotations from each camera.
CFC. The Caltech Fish Counting Dataset (CFC) [30] is a domain generalization benchmark sourced from fisheries monitoring, where sonar video is used to detect and count migrating salmon. The detection task consists of a single class (“fish”) and domain shift is caused by real-world environmental differences between camera deployments. We identify this application as an opportunity to study the generality of DAOD methods due to its stark differences with existing DAOD benchmarks—specifically, sonar imagery is grayscale, has low signal-to-noise ratios, and foreground objects are difficult to distinguish from the background—however CFC focuses on generalization rather than adaptation and does not include the data needed for DAOD.
CFC-DAOD We introduce an extension to CFC, deemed CFC-DAOD, to enable the study of DAOD in this application domain. The task is to adapt from a source location—“Kenai”, i.e. the default training set from CFC—to a difficult target location, “Channel”. We collected an additional 168k bounding box annotations in 29k frames sampled from 150 new videos captured over two days from 3 different sensors on the “Channel” river (see Fig. 3). For consistency, we closely followed the video sampling protocol used to collect the original CFC dataset as described by the authors (see [30]). Our addition to CFC is crucial for DAOD as it adds an unsupervised training set for domain adaptation methods and a supervised training set to train oracle methods. We keep the original supervised Kenai training set from CFC (132k annotations in 70k images) and the original Channel test set (42k annotations in 13k images). We note this is substantially larger than existing DAOD benchmarks (CS contains 32k instances in 3.5k images, and Sim10k contains 58k instances in 10k images). See the supplemental material for more dataset statistics. We make the dataset public.
6 Experiments
In this section we propose an updated benchmarking protocol for DAOD (Sec. 6.1) that allows us to fairly analyze the performance of ALDI++ compared to prior work (Sec. 6.2) and conduct extensive ablation studies (Sec. 6.3).
Datasets. We perform experiments on Cityscapes $\rightarrow$ Foggy Cityscapes, Sim10k $\rightarrow$ Cityscapes, and CFC Kenai $\rightarrow$ Channel. In addition to being consistent with prior work, these datasets represent three common adaptation scenarios capturing a range of real-world challenges: weather adaptation, Sim2Real, and environmental adaptation, respectively. We note that there have been inconsistencies in prior work in terms of which ground truth labels for Cityscapes are used. We use the Detectron2 version.
Metrics. For all experiments we report the PascalVOC metric of mean Average Precision with IoU $\geq 0.5$ (“AP50”) [15]. This is consistent with prior work on Cityscapes, Foggy Cityscapes, Sim10k, and CFC.
6.1 A New Benchmarking Protocol for DAOD
Figure 4: Revisiting source-only and oracle models in DAOD. We argue that in order to provide a fair measure of domain adaptation performance in DAOD, source-only and oracle models must utilize the same non-adaptive architectural and training components as methods being studied. In the case of Align and Distill-based approaches, this means source-only and oracle models must have access to the same set of source augmentations and EMA as DAOD methods. We see that these upgrades significantly improve source-only performance on target-domain data (+7.2 AP50 on Foggy Cityscapes), even though the source-only model has never seen any target-domain data, and these upgrades also improve oracle performance. Overall, these results set more challenging and realistic performance targets for DAOD methods.
Revisiting source-only and oracle models. Here we address P1: Improperly constructed source-only and oracle models, leading to overestimation of performance gains. The goal of DAOD is to develop adaptation techniques that use unlabeled target-domain data to improve target-domain performance. Thus, in order to properly isolate adaptation-specific techniques, any technique that does not need target-domain data to run should also be used by source-only and oracle models. In our case, this means that source-only and oracle models should also utilize the same strong augmentations and EMA updates as DAOD methods. In Fig. 4 we illustrate the resulting source-only and oracle models, and show that including these components significantly improves both source-only and oracle model performance (+7.2 and +2.6 AP50 on Foggy Cityscapes, respectively). This has significant implications for DAOD research: because source-only and oracle models have not been constructed with equivalent components, performance gains stemming from better generalization have until now been misattributed to DAOD. With properly constructed source-only and oracle models, the gains from DAOD are much more modest (see Fig. 5).
Modernizing architectures. We next address P3: (b) Outdated model architectures leading to overestimation of methods’ generality. Prior art in DAOD has used older backbones (e.g. VGG-16) in order to compare to previously-published results. To investigate whether conclusions drawn from these results generalize to modern experimental settings, our experiments utilize a modern detection framework [54] with default settings including multi-scale input transforms and COCO pre-training. We use a ResNet-50 backbone [23] with Feature Pyramid Network [35] and ViTDet [33]. We provide more details in the supplementary. Source-only and oracle models also receive these upgrades.
6.2 Fair Comparison and State-of-the-Art Results
We compare ALDI++ with reimplementations of five state-of-the-art DAOD methods on top of our framework: UMT [12], SADA [10], PT [8], MIC [26], and AT [34]; see Tab. 1 for the ALDI settings used to reproduce them. We use the fair benchmarking protocol proposed in Sec. 6.1. Results are shown in Fig. 5. All methods (including ALDI++) use the same settings for all benchmarks.
ALDI++ is state-of-the-art on CS normal-$\rightarrow$ FCS, Sim10k normal-$\rightarrow$ CS, and CFC Kenai normal-$\rightarrow$ Channel. ALDI++ outperforms the previous state-of-the-art by +3.5 AP50 on CS $\rightarrow$ FCS, +5.7 AP50 on Sim10k $\rightarrow$ CS (where ours is the only method to outperform a fair source-only model), and +2.0 AP50 on CFC Kenai $\rightarrow$ Channel. Further, we achieve near-oracle level performance on CS $\rightarrow$ FCS and CFC Kenai $\rightarrow$ Channel (0.4 and 0.9 AP50 away, respectively), while other methods close less than half the gap between source-only and oracle models.
Modern architectures and fair source-only models create a paradigm shift for DAOD benchmarking. Our benchmarking protocol provides a dramatic reset for DAOD performance bounds. A fair and modern source-only model—trained without ever seeing any target domain data—achieves a higher target AP50 than all previously-published DAOD methods (see Fig. 1). Similarly, we see a 57% increase in oracle performance compared to prior work.
Figure 5: ALDI++ outperforms all prior work on CS normal-$\rightarrow$ FCS, Sim10k normal-$\rightarrow$ CS, and CFC Kenai normal-$\rightarrow$ Channel. We provide a fair comparison of ALDI++ with five existing state-of-the-art approaches by upgrading them to use the ALDI framework. We see that some prior methods continue to provide benefit on top of a modern architecture but others lag behind modern source-only models. Our method ALDI++ outperforms prior work on all datasets studied by a significant margin: +3.5 AP50 on CS normal-$\rightarrow$ FCS, +5.7 AP50 on Sim10k normal-$\rightarrow$ CS, and +2.0 AP50 on CFC Kenai normal-$\rightarrow$ Channel. Notably, ALDI++ is the only method to outperform a source-only model on Sim10k normal-$\rightarrow$ CS.
Relative performance of all methods decreases compared to fair source-only and oracle models. Re-implementing SOTA methods in ALDI improves absolute performance of all methods; however performance decreases compared to source-only models. There are several instances where modernized DAOD methods are actually worse than a fair source-only model. Notably, a source-only model outperforms upgraded versions all previously-published work on Sim10 $\rightarrow$ CS. We also see that no state-of-the-art methods outperform a fair oracle on any dataset, in contrast to claims made by prior work [34, 8, 5].
The ranking of methods differs across datasets and architectures. MIC and AT are consistently the top-performing methods across all datasets. UMT exhibits variable performance due to the differences in the difficulty of image generation across datasets (see supplemental for examples). SADA underperforms other methods on CS $\rightarrow$ FCS and CFC Kenai $\rightarrow$ Channel, but closes this gap on the more difficult Sim10k $\rightarrow$ CS. These differences highlight the importance of benchmarking in a modern context, as we see that previously-published methods are not always complementary to general advancements in object detection. These results also demonstrate the utility of CFC-DAOD as another point of comparison for DAOD methods.
ALDI++ is compatible with new detector variants. We upgrade ALDI++ to use VitDet [33]. Since VitDet is a two-stage architecture based on Faster R-CNN, this requires no modifications to our multi-task distillation loss. We show that ALDI++ continue to show improvements over an upgraded VitDet source-only model (see Fig. 1 for CS $\rightarrow$ FCS and supplemental for other datasets). We see there is a larger gap between the ViT ALDI++ and the ViT oracle, indicating the potential for future work to improve performance.
6.3 Ablation Studies
In this section we ablate the performance of each component of ALDI on CS $\rightarrow$ FCS. For each ablation, unless otherwise specified we begin with the settings shown in Tab. 2, ablating one column at a time. Our default burn-in uses the same augmentations as the Fixed strategy from Tab. 1 but uses early stopping for model selection, and all self-training runs are initialized with the same burned-in checkpoint for fair comparison. See Tab. 1 for additional definitions.
Table 2: Starting point for all ablation experiments.
Method $\theta_{stu},\theta_{tch}$ (Burn-in) $T_{src}$ $T_{tgt}$ $\frac{B_{tgt}}{B}$ Postprocess $L_{distill}$ $L_{align}$
Ablation start Weak augs, early stopping F, M F, M, J, C 0.5 Thresh, NMS Hard
$\theta_{stu}$, $\theta_{tch}$ Network initialization (burn-in). In Fig. 6a we analyze the effects of our proposed burn-in strategy (see Sec. 4). We measure performance in terms of target-domain AP50 as well as convergence time, defined as the training time at which the model first exceeds 95% of its final target-domain performance. We compare our approach with: (1) No dataset-specific burn-in, i.e. starting with COCO weights, and (2) The approach used by past work—using a fixed burn-in duration, e.g. 10k iterations. We find that our method results in significant improvements in both training speed and accuracy, leading to upwards of 10% improvements in AP50 and reducing training time by a factor of 10 compared to training without burn-in.
[…]
$T_{src}$ Source augmentations. In Fig. 6b we ablate the set of source-domain data augmentations. We compare using weak augmentations (random flipping and multi-scale training), strong augmentations (color jitter and random erasing), and a combination of weak and strong, noting that prior works differ in this regard but do not typically report the settings used (see Tab. 1). We find that using strong source augmentations on the entire source-domain training batch outperforms weak augmentations and a combination of both.
$T_{tgt}$ Target augmentations. In LABEL:tab:target_aug we investigate the use of different augmentations for target-domain inputs to the student model. (We note that weak augmentations are always used for target-domain inputs to the teacher in accordance with prior work). We see that stronger augmentations consistently improve performance, with best performance coming from the recently-proposed MIC augmentation [26].
Table 3: (a) Effects of target-domain augmentation on self-training. Augmentations applied to student inputs ($T_{tgt}$ in Fig. 2). Stronger augmentations improve performance considerably. (b) Effects of distillation objectives on self-training. We compare hard targets—used by most prior art, which thresholds teacher predictions to create pseudo-labels—with our proposed soft targets. Soft targets can improve overall performance. Results are the mean and 95% confidence intervals over 3 runs. (c) Feature alignment has diminishing returns. Alignment objectives $L_{align}$ without self-training provides up to 2.6 AP50 of benefit (left column), but diminishes to 0.2 AP50 additional gain when used alongside self-training (right column).
[…]
Source-only model 51.9
Weak (scale & flip) 52.6
+ Color jitter 59.0
+ Color jitter + Erase 63.1
+ Color jitter + MIC 64.3
[…]
Source-only 59.1
Hard targets 63.7 $\pm$ 0.1
Soft targets 64.0 $\pm$ 0.4
[…]
Source-only 59.1
61.7 (+2.6)
63.7 (+4.6)
63.9 (+4.8)
[…]
$L_{distill}$ Self-distillation. In LABEL:tab:distill_ablate we analyze the effects of our proposed multi-task soft distillation approach (see Sec. 4). Note that for these experiments, the starting model is ALDI++ rather than the simple model in Tab. 2. We compare our approach with the “hard” pseudo-label approach used by prior work, where teacher predictions are post-processed with non-maximum suppression and a hard confidence threshold of 0.8 [26, 34, 12, 37]. For our proposed “soft” distillation method, we first sharpen teacher predictions at both detector stages using a sigmoid for objectness predictions and a softmax for classification predictions, both with a default temperature of 1. We see that our proposed soft targets improve performance compared to hard targets.
[…]
7 Discussion and Conclusions
In this work we proposed the ALDI framework and an improved DAOD benchmarking methodology, providing a critical reset for the DAOD research community; a new dataset CFC-DAOD, increasing the diversity and real-world applicability of DAOD benchmarks; and a new method ALDI++ that advances the state-of-the-art. We conclude with key findings.
Network initialization has an outsized impact. We find that general advancements in computer vision eclipse progress in DAOD: a Resnet50-FPN source-only model outperforms all VGG-based DAOD methods, and a VitDet source-only model outperforms all Resnet50-FPN based DAOD methods. Similarly, simply adding stronger augmentations and EMA to source-only models leads to better target-domain performance than some adaptation methods, and including these upgrades during network initialization (burn-in) improves adaptation performance as well.
DAOD techniques are helpful, but do not achieve oracle-level performance as previously claimed [34, 8, 5]. Top-performing DAOD methods, including ALDI++, demonstrate improvements over source-only models (see Fig. 1 and Fig. 5). However, in contrast to previously-published results, no DAOD method reaches oracle-level performance, suggesting there is still room for improvement. The gap between DAOD methods and oracles is even larger for stronger architectures like VitDet. This is a promising area for future research.
Benchmarks sourced from real-world domain adaptation challenges can help the community develop generally useful methods. We find that DAOD methods do not necessarily perform equivalently across datasets (see Fig. 5). Diverse benchmarks are useful to make sure we are not overfitting to the challenges of one particular use case, while exposing and supporting progress in impactful applications. Our contributed codebase and benchmark dataset provide the necessary starting point to enable this effort.
A lack of transparent comparisons has incentivized incremental progress in DAOD. Most highly-performant prior works in DAOD are some combination of DANN [17] (2016) and Mean Teacher [50] (2017) plus custom training techniques. Without fair comparisons it has been possible to propose near-duplicate methods that still achieve “state-of-the-art” performance due to hyperparameter tweaks. Our method ALDI++ establishes a strong point of comparison for Align and Distill-based approaches that will require algorithmic innovation to surpass.
Validation is the elephant in the room. All of our experiments, and all previously published work in DAOD, utilize a target-domain validation set to perform model and hyperparameter selection. This violates a key assumption in unsupervised domain adaptation: that no target-domain labels are available to begin with. Prior work has shown that it may not be possible to achieve performance improvements in domain adaptation at all under realistic validation conditions [38, 39, 31]. Therefore our results (as well as previously-published work) can really only be seen as an upper bound on DAOD performance. While this is valuable, further research is needed to develop effective unsupervised validation procedures for DAOD.
Acknowledgements. This material is based upon work supported by: NSF CISE Graduate Fellowships Grant #2313998, MIT EECS department fellowship #4000184939, MIT J-WAFS seed grant #2040131, and Caltech Resnick Sustainability Institute Impact Grant “Continuous, accurate and cost-effective counting of migrating salmon for conservation and fishery management in the Pacific Northwest.” Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NSF, MIT, J-WAFS, Caltech, or RSI. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources [45]. We also thank the Alaska Department of Fish and Game for their ongoing collaboration and for providing data, and Sam Heinrich, Neha Hulkund, Kai Van Brunt, and Rangel Daroya for helpful feedback.